OpenStudy (anonymous):

solve limit(n+1)^1/ln(n+1)

OpenStudy (fibonaccichick666):

to start, you can't solve a limit that doesn't have a limit

OpenStudy (anonymous):

Let's rewrite \(\ln(n+1)=u\) so \(n+1=e^u\). Initially, assuming you had \(n\to\infty\), we now consider \(u=\ln(n+1)\to\infty\):$$\lim_{u\to\infty}(e^u)^{1/u}=\lim_{u\to\infty}e^{u/u}=\lim_{u\to\infty}e=e$$

OpenStudy (anonymous):

We simply replaced \(\ln(n+1)\) with \(u\) and so we had to replace \(n+1=e^u\). It simplifies dramatically!